OpenStudy (anonymous):

can someone plz help me this is my last day of school and i need help i will medal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Esmeralda and Heinz are working to graph a polynomial function, f(x). Esmeralda says that the third degree polynomial has four intercepts. Heinz argues that the function only crosses the x axis three times Is there a way for both of them to be correct? Explain your answer

OpenStudy (anonymous):

"Esmeralda says that the third-degree polynomial has four intercepts." A third-degree polynomial has at the most three real roots and so it can cross the x-axis at the most three times. So max. x-intercepts is 3. But here it says four intercepts which could include BOTH x and y intercepts. So both could be correct. IF Esmeralda said: "the third-degree polynomial has ** four X intercepts **" then Esmeralda will be wrong and Heinz will be correct.

OpenStudy (anonymous):

oh ok that makes sense do you think that you could help me with one more?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x) \[g(x)=x^3-x^2-4x+4\]\[g(x)=x^3+2x^2-9x-18\]\[g(x)=x^3-3x2-4x+12\]\[g(x)=x^3+2x^2-25x-50\]\[g(x)=2x^3+14x^2-2x-14\]

OpenStudy (anonymous):

@SneliS

OpenStudy (anonymous):

Take the first function g(x)=x 3 –x 2 –4x+4 Graph it and find where it crosses the x-axis and those will be the roots or zeros of g(x). You can use this online graphing website. Type x^3 – x^2 – 4x + 4 on the left column and it will plot it for you.

OpenStudy (anonymous):

use this https://www.desmos.com/calculator

OpenStudy (anonymous):

ok thx but what does it mean about the other key features?

OpenStudy (anonymous):

hold on did u use the calculator if so what did u get?

OpenStudy (anonymous):

(-2,0) (1,0) and (2,0)

OpenStudy (anonymous):

@Hero @Preetha @Michele_Laino @Abhisar @Australopithecus can u plz help me with this last part

OpenStudy (anonymous):

f(x) means function of x. The value of f depends on the value of x. f(x) = x + 2 Here, the left hand side f depends on the value of x. If x = 1, f(1) = 3. If x =2, f(2) = 4, If x = 5, f(5) = 7. g(x) is just another function of x. When plotting the f(x) or g(x) will be along the y-axis and the x will be along the x axis.

OpenStudy (anonymous):

ok but what are the key features of g(x) @SneliS

OpenStudy (anonymous):

@SneliS

OpenStudy (anonymous):

idk?? honestly

OpenStudy (anonymous):

oh ok thats cool thx anyway tho